# From Dormant to Deleted: Tamper-Resistant Unlearning Through Weight-Space Regularization

### v1

@@ 6.1 A New Class of Tamper-Resistant Unlearning Methods
Connectivity-Based Fine-Tuning (CBFT). We employ the method from (lubana2023mechanistic, ) (originally proposed to obtain models focusing on distinct recognition mechanisms). We maximize the loss on the midpoint between the pretrained model and the current unlearned model on examples from the retain set as well as the forget set, while only minimizing the loss on the retain set for the unlearned model. This attempts to add a high-loss barrier in between, while still retaining model utility on the retain set for the final unlearned model. We use a small weighting factor of 0.001 on the loss maximization term. Similar to (lubana2023mechanistic, ), we ignore the loss maximization term if the loss magnitude exceeds a value of 50.
@@
Findings. Fig. 2 shows that Weight Distortion and Weight Dist Reg are significantly more tamper-resistant in comparison to prior approaches. However, we observe that CBFT is less effective than Weight Dist Reg and Weight Distortion across the board. We hypothesize that this is due to acting on model outputs and only indirectly influencing the weight-space. Indeed, while CBFT does create a larger loss barrier than other methods (Fig. 4), the $L_{2}$ norm between the pretrained and unlearned parameters is relatively low (Fig. 5). Overall, out of the two weight-space diagnostic tools, we find that the $L_{2}$ norm of the difference in model parameters is more reliable in predicting tamper-resistance. Fig. 3 shows that both Weight Dist Reg and CBFT, when applied as a second phase on an initial safeguard, can substantially improve its tamper-resistance, unlike TAR. The only exception is if the initial safeguard is Weight Distortion, which already has sufficient tamper-resistance.
7 Discussion and Conclusion
@@
On the trade-off between tamper-resistance and test accuracy. As discussed previously, there are inherent trade-offs in unlearning, between forgetting the specified examples while maintaining utility (measured via test accuracy) (kurmanji2024scrub, ; deeb2024unlearning, ; triantafillou2024unlearningcomp, ). Here, we discover a different trade-off between resisting relearning attacks, and test accuracy. Indeed, Fig. 2 shows that the methods that are best at defending against these attacks are the ones with the lowest test accuracy. Having surfaced this fundamental tension, we hope future work improves on the current Pareto frontier formed by our new methods.
@@
Findings hold across datasets/architectures. The results presented in the paper are consistent across models i.e., ResNet-34 on CIFAR-10 as presented in Fig. 9Appendix D. Furthermore, these results are also consistent across different datasets, i.e., CIFAR-100, as highlighted in Fig. 8Appendix C.
@@
The role of the retain set for relearning. Given the surprising finding that we can recover forget set accuracy while fine-tuning on only the retain set, we ask: are the retain set examples necessary for this to occur, or could we have used any other examples from the same distribution? We investigate this in Appendix F where we replace the retain set with different sets, and fine-tune the unlearned model on those different sets (combined with the ‘relearning set’, which is a subset of the forget set, as before). Notably, replacing the retain set with test examples (that the model was not previously trained on) causes the relearning effect to be a bit less pronounced, especially when 0, or few, relearning examples are used, highlighting the importance of using training rather than held-out data for inducing relearning. This observation relates to other recent findings on anticipatory knowledge reawakening when exposing the model to a repeated sequence of documents. In this scenario, as the model processes documents in a fixed order, it unexpectedly begins to recover an increasing amount of information about a previously seen example even before encountering that example again (yang2024reawakening, ).
@@
Summary and take-aways. We showed that unlearning methods are susceptible to relearning attacks where the forget set accuracy can be recovered simply by fine-tuning on the retain set. For atypical examples in particular, there is a stark contrast between the relearning patterns of unlearning methods compared to retraining from scratch. Based on weight-space analysis, we suggest two diagnostic tools for understanding tamper-resistance, and propose simple methods that yield state-of-the art tamper-resistant results, revealing new pathways for improving unlearning. The authors of  (sepahvand2025selective, ) argued that unlearning algorithms that operate on representations rather than at the level of outputs may be more robust at defending some types of attacks. Our findings take this discussion further, showing that unlearning methods that operate at either the level of the model outputs or the representations, without any constraint on model weights (which include methods such as SCRUB, Circuit Breakers, and Gradient Ascent, etc) should struggle to induce robustness against relearning attacks. On the other hand, methods that directly or indirectly push the pretrained and the unlearned models apart via any intervention (which includes distortion model weights, regularizing the model to decay parameter magnitude, or even directly pushing the models apart using an explicit loss term) should be significantly more robust to relearning attacks. Our proposed methods for increasing tamper-resistance exemplify this and we hope future work builds on these further.
@@ Acknowledgements
The authors would like to acknowledge useful discussions with Yanzhi Chen and Ilia Shumailov regarding unlearning, and susceptibility to relearning. AW acknowledges support from Turing AI Fellowship under grant EP/V025279/1, the Alan Turing Institute, and the Leverhulme Trust via CFI.
@@ Appendix A Tamper-Resistance with Typical Examples for the Forget Set
As highlighted in Section 5, typical examples can be particularly easy to unlearn as simply ignoring these examples during training (i.e., a no-op) yields similar predictions as the retrained from scratch model [19]. This is because they are, by definition, easy to predict regardless of whether they are part of the training set or not  [9, 20, 1, 33]. Generally, because typical examples are more likely to be predicted correctly, they are easier to relearn as well. For the sake of completeness, we include the results for typical examples in Fig. 6. As evident from the figure, even the retrained from scratch model already achieves perfect accuracy on the forget set, highlighting that unlearning, and in turn, relearning are both trivial for typical examples. Therefore, we primarily focused on atypical examples in Section 5. We see that methods such as Random Relabeling, while completing forgetting the forget set, deviates in a very significant way from the retrain from scratch baseline, which is what we compare against.
@@ Appendix B Tamper-Resistance with Random Set of Examples for the Forget Set
We further evaluate performance in the case of random selection of examples, rather than just typical examples as evaluated in Appendix A. We visualize these results in Fig. 7. Since randomly selected examples are predominantly typical [9, 8, 1, 33], retrain from scratch already achieves near-perfect accuracy. Furthermore, we observe only minor differences between methods. Consequently, we focused on atypical examples for the forget set in the main paper (Section 5) as these examples are hard to predict without being part of the training set, marking clearly the impact of relearning attacks which goes beyond model generalization.
@@ Appendix C Tamper-Resistance on a More Complex Dataset (CIFAR-100)
Figure 8: Comparison between test set accuracy and accuracy on the held-out part of the forget set $\mathcal{D}_{F_{ho}}$ on CIFAR-100 and ResNet-18, where the forget set is comprised of atypical examples from the ‘apple’ class. We see significant variation among different methods on the more complex CIFAR-100 dataset, which is comparable to the results we observed with class-agnostic unlearning on CIFAR-10 in Fig. 2 (bottom).
We predominantly focus on CIFAR-10 dataset in the main paper (Section 5). We evaluate the susceptibility on a more complex CIFAR-100 dataset in Fig. 8 in order to understand the impact of the dataset. We only look at sub-class unlearning in this case, focusing on the ‘apple’ class for unlearning. Due to the higher complexity of the dataset, we observe lower test accuracies as evident from the scale of the x-axis. Retrain from scratch baseline achieves a low forget set accuracy. Furthermore, we see a wider spread of different methods on this more complex dataset, with the trend looking similar to the class-agnostic results presented in Fig. 2 (as it provided a more complex forget set comprising of the most atypical examples from the dataset). It is interesting to note that CBFT [26], which was not tamper-resistant on CIFAR-10, demonstrated significant resistance on CIFAR-100.
Appendix D Tamper-Resistance of a Larger Model
@@
All our prior results focused on the smaller model, i.e., ResNet-18 [16]. In order to understand the impact of model size, we evaluate the susceptibility against relearning attacks on ResNet-34 (with almost double the number of parameters compared to ResNet-18, going from $\sim 11M$ to $\sim 21M$). The results are visualized in Fig. 9. Note that we directly transfer the hyperparameter settings from our ResNet-18 experiments. Hence, we found that Weight Dist Reg, which was the most competitive method in terms of robustness, became susceptible to relearning. It is worth noting that the test set accuracy as well as the forget set accuracy for Weight Dist Reg in this case deviates significantly from all prior results, where it demonstrated consistently lower accuracies, highlighting a potential deficiency of the selected hyperparameters. Catastrophic Forgetting and Weight Dist Reg exhibit high robustness against relearning, which are instantiations of our framework.
Appendix E Evolution of Forget Set Accuracy During Relearning
Figure 10: Comparison between relearning time and accuracy on the held-out part of the forget set $\mathcal{D}_{F_{ho}}$ on CIFAR-10 and ResNet-18, where the forget set is comprised of atypical examples from the ‘airplane’ class. The figure indicates that many methods achieve near-perfect recovery of unlearned knowledge with only a small amount of model fine-tuning, without even assuming access to the unlearned examples. This figure is an extension of the results presented in Fig. 2 (top).
Figure 11: Comparison between test set accuracy and accuracy on the held-out part of the forget set $\mathcal{D}_{F_{ho}}$ on CIFAR-10 and ResNet-18, where the forget set is comprised of atypical examples from all classes. This figure is an extension of the results presented in Fig. 2 (bottom). We find that atypical examples, when selected from all the classes, are significantly harder in comparison to examples selected from a particular class, and hence, achieve a better separation between different methods in comparison to sub-class unlearning in Fig. 2 (top).
In order to complement the results in Fig. 2 (top) and Fig. 3, we visualize the forget set accuracy evolution in Fig. 10. We see that methods that are most stable are nearly unaffected by increasing training time, highlighting that more training time might not be sufficient to increase their susceptibility further.
Similarly, Fig. 11 complements the results on class-agnostic unlearning in Fig. 2 (bottom), where we visualize both the line plots as well as the results from two-phase training strategies. We visually observe best separation between different methods on the line plot in this case of class-agnostic unlearning.
Appendix F Relearning with examples from the same distribution, which are distinct from the retain set
@@
The results are presented in Fig. 12. We observe that using the retain set $\mathcal{D}_{R}$ achieves higher accuracy on the forget set compared to other choices. However, these differences diminish as the number of relearning examples increases, since these examples directly represent the unlearned knowledge and may become the dominant factor in recovery. When the test set is used for relearning (instead of the retain set), we unsurprisingly observe perfect accuracy on the test set, as it serves as the training set in this scenario. These results indicate that relearning is more effective when using examples that were seen during training, rather than a held-out set of examples.
@@
Appendix G Computational Resources

### v2

@@ 6.1 A New Class of Tamper-Resistant Unlearning Methods
Connectivity-Based Fine-Tuning (CBFT). We employ the method from (lubana2023mechanistic, ), which was originally proposed to obtain models focusing on distinct recognition mechanisms. We maximize the loss on the midpoint between the pretrained model and the current unlearned model on examples from the retain set as well as the forget set, while only minimizing the loss on the retain set for the unlearned model. This attempts to add a high-loss barrier in between, while still retaining model utility on the retain set for the final unlearned model. We use a small weighting factor of 0.001 on the loss maximization term. Similar to (lubana2023mechanistic, ), we ignore the loss maximization term if the loss magnitude exceeds a value of 50.
@@
Findings. Fig.˜2 shows that Weight Distortion and Weight Dist Reg are significantly more tamper-resistant in comparison to prior approaches. However, we observe that CBFT is less effective than Weight Dist Reg and Weight Distortion across the board. We hypothesize that this is due to acting on model outputs and only indirectly influencing the weight-space. Indeed, while CBFT does create a larger loss barrier than other methods (Fig.˜4), the $L_{2}$ norm between the pretrained and unlearned parameters is relatively low (Fig.˜5). Overall, out of the two weight-space diagnostic tools, we find that the $L_{2}$ norm of the difference in model parameters is more reliable in predicting tamper-resistance. Fig.˜3 shows that both Weight Dist Reg and CBFT, when applied as a second phase on an initial safeguard, can substantially improve its tamper-resistance, unlike TAR. The only exception is if the initial safeguard is Weight Distortion, which already has sufficient tamper-resistance.
7 Analysis and Discussion
@@
On the trade-off between tamper-resistance and test accuracy. As discussed previously, there are inherent trade-offs in unlearning, between forgetting the specified examples while maintaining utility (measured via test accuracy) (kurmanji2024scrub, ; deeb2024unlearning, ; triantafillou2024unlearningcomp, ). Here, we discover a different trade-off between resisting relearning attacks, and test accuracy. Indeed, Fig.˜2 shows that the methods that are best at defending against these attacks are the ones with the lowest test accuracy. Having surfaced this fundamental tension, we hope future work improves on the current Pareto frontier formed by our new methods.
@@
Findings hold across datasets/architectures. The results presented in the paper are consistent across models i.e., ResNet-34 on CIFAR-10 as presented in Fig.˜10Appendix˜E. Furthermore, these results are also consistent across different datasets, i.e., CIFAR-100, as highlighted in Fig.˜9Appendix˜C. Finally, our findings are also consistent on a higher-resolution Imagenette dataset (imagenette, ) as highlighted in Table˜1Appendix˜D.
Efficiency of Unlearning Algorithms. We used a fix budget of 100 epochs for unlearning to avoid confounding from differences in optimization efficiency. In order to justify the use of approximate unlearning algorithms instead of exact unlearning (i.e., retrain from scratch) with a training budget comparable to pretraining, we evaluate the efficiency of unlearning algorithms in Table˜3Appendix˜I. The results show that our methods are much more efficient at unlearning (and consequently resisting re-emergence of knowledge) even after a single epoch of unlearning compared to other approaches.
Difference between Quantization and Relearning Attacks. Language models have been shown to be susceptible to quantization attacks (zhang2024catastrophic, ), where quantization after unlearning is sufficient to recover unlearned knowledge from the model. We evaluate susceptibility to quantization in Fig.˜6. Interestingly, our vision models are not vulnerable to these quantization attacks: forget set accuracy stays at the same level as the full-precision unlearned model regardless of the number of bits until it gets close to 8 bits, at which point the model collapses. Thus, the spontaneous-recovery phenomenon we explore via relearning attacks is distinct from the recovery obtained via quantization. We hypothesize that the primary reason why quantization attacks succeed in language models is due to their larger number of parameters (on the order of 100 billion parameters (dubey2024llama, ; liu2024deepseek, )), resulting in very small changes in weights during unlearning. Hence, models can easily revert to their previous values with quantization. This problem is further exacerbated with models that are trained with quantization in mind. Such models can immediately latch onto the previous values that were regularized to be easily quantizable (jung2019learning, ). Similar to our findings, (zhang2024catastrophic, ) hypothesized that small differences are the cause of this susceptibility to quantization attacks. They further proposed increasing the learning rate as a potential mechanism to induce larger differences between the pretrained and unlearned models (zhang2024catastrophic, ), while our work introduces more principled approaches for the same purpose.
Figure 6: Comparison between the number of quantized bits and accuracy of the model for different unlearning methods on CIFAR-10 and ResNet-18, where the forget set is comprised of atypical examples. In contrast to language models, where quantization has been shown to be responsible for the re-emergence of unlearned knowledge (zhang2024catastrophic, ), we observe that quantization has little to no impact on forget set accuracy until the point where full performance is recovered. Note that the accuracy of 100% at the start in most cases is an artifact of the model prediction collapsing to just class 0 (whose instances have been selected as our forget set).
@@
The role of the retain set for relearning. Given the surprising finding that we can recover forget set accuracy while fine-tuning on only the retain set, we ask: are the retain set examples necessary for this to occur, or could we have used any other examples from the same distribution? We investigate this in Appendix˜G where we replace the retain set with different sets, and fine-tune the unlearned model on those different sets (combined with the ‘relearning set’, which is a subset of the forget set, as before). Notably, replacing the retain set with test examples (that the model was not previously trained on) causes the relearning effect to be a bit less pronounced, especially when 0, or few, relearning examples are used, highlighting the importance of using training rather than held-out data for inducing relearning. This observation relates to other recent findings on anticipatory knowledge reawakening when exposing the model to a repeated sequence of documents. In this scenario, as the model processes documents in a fixed order, it unexpectedly begins to recover an increasing amount of information about a previously seen example even before encountering that example again (yang2024reawakening, ).
8 Conclusion
@@
We showed that unlearning methods are susceptible to relearning attacks where the forget set accuracy can be recovered simply by fine-tuning on the retain set. For atypical examples in particular, there is a stark contrast between the relearning patterns of unlearning methods compared to retraining from scratch. Based on weight-space analysis, we suggest two diagnostic tools for understanding tamper-resistance, and propose simple methods that yield state-of-the art tamper-resistant results, revealing new pathways for improving unlearning. The authors of  (sepahvand2025selective, ) argued that unlearning algorithms that operate on representations rather than at the level of outputs may be more robust at defending some types of attacks. Our findings take this discussion further, showing that unlearning methods that operate at either the level of the model outputs or the representations, without any constraint on model weights (which include methods such as SCRUB, Circuit Breakers, and Gradient Ascent, etc) should struggle to induce robustness against relearning attacks. On the other hand, methods that directly or indirectly push the pretrained and the unlearned models apart via any intervention (which includes distorting model weights, regularizing the model to decay parameter magnitude, or even directly pushing the models apart using an explicit loss term) should be significantly more robust to relearning attacks. Our proposed methods for increasing tamper-resistance exemplify this and we hope future work builds on these further.
@@ Acknowledgements
The authors would like to acknowledge useful discussions with Yanzhi Chen and Ilia Shumailov regarding unlearning, and susceptibility to relearning. We are also very thankful to the anonymous reviewers for their useful feedback on the initial submission. AW acknowledges support from Turing AI Fellowship under grant EP/V025279/1, the Alan Turing Institute, and the Leverhulme Trust via CFI.
@@ Appendix A Tamper-Resistance with Typical Examples for the Forget Set
As highlighted in Section˜5, typical examples can be particularly easy to unlearn as simply ignoring these examples during training (i.e., a no-op) yields similar predictions as the retrained from scratch model [21]. This is because they are, by definition, easy to predict regardless of whether they are part of the training set or not  [10, 22, 1, 39]. Generally, because typical examples are more likely to be predicted correctly, they are easier to relearn as well. For the sake of completeness, we include the results for typical examples in Fig.˜7. As evident from the figure, even the retrained from scratch model already achieves perfect accuracy on the forget set, highlighting that unlearning, and in turn, relearning are both trivial for typical examples. Therefore, we primarily focused on atypical examples in Section˜5. We see that methods such as Random Relabeling, while completing forgetting the forget set, deviates in a very significant way from the retrain from scratch baseline, which is what we compare against.
@@ Appendix B Tamper-Resistance with Random Set of Examples for the Forget Set
We further evaluate performance in the case of random selection of examples, rather than just typical examples as evaluated in Appendix˜A. We visualize these results in Fig.˜8. Since randomly selected examples are predominantly typical [10, 9, 1, 39], retrain from scratch already achieves near-perfect accuracy. Furthermore, we observe only minor differences between methods. Consequently, we focused on atypical examples for the forget set in the main paper (Section˜5) as these examples are hard to predict without being part of the training set, marking clearly the impact of relearning attacks which goes beyond model generalization.
@@ Appendix C Tamper-Resistance on a More Complex Dataset (CIFAR-100)
Figure 9: Comparison between test set accuracy and accuracy on the held-out part of the forget set $\mathcal{D}_{F_{ho}}$ on CIFAR-100 and ResNet-18, where the forget set is comprised of atypical examples from the ‘apple’ class. We see significant variation among different methods on the more complex CIFAR-100 dataset, which is comparable to the results we observed with class-agnostic unlearning on CIFAR-10 in Fig.˜2 (bottom).
We predominantly focus on CIFAR-10 dataset in the main paper (Section˜5). We evaluate the susceptibility on a more complex CIFAR-100 dataset in Fig.˜9 in order to understand the impact of the dataset. We only look at sub-class unlearning in this case, focusing on the ‘apple’ class for unlearning. Due to the higher complexity of the dataset, we observe lower test accuracies as evident from the scale of the x-axis. Retrain from scratch baseline achieves a low forget set accuracy. Furthermore, we see a wider spread of different methods on this more complex dataset, with the trend looking similar to the class-agnostic results presented in Fig.˜2 (as it provided a more complex forget set comprising of the most atypical examples from the dataset). It is interesting to note that CBFT [30], which was not tamper-resistant on CIFAR-10, demonstrated significant resistance on CIFAR-100.
Appendix D Tamper-Resistance on a Higher-Resolution Dataset (Imagenette)
Model Label Test Acc (%) Forget Set Acc (%)
Unlearned Retrain from Scratch 87.34% 67.53%
Unlearned SCRUB 85.81% 85.71%
Unlearned Weight Distortion 83.16% 74.03%
Unlearned Weight Dist Reg 83.72% 84.42%
0 Relearn Retrain from Scratch 86.40% 63.53%
0 Relearn SCRUB 86.47% 93.40%
0 Relearn Weight Distortion 84.71% 78.86%
0 Relearn Weight Dist Reg 84.31% 73.87%
10 Relearn Retrain from Scratch 86.73% 64.16%
10 Relearn SCRUB 86.65% 94.21%
10 Relearn Weight Distortion 84.62% 81.09%
10 Relearn Weight Dist Reg 84.83% 75.77%
Table 1: Test and forget set accuracy across unlearning and relearning conditions when evaluating on the higher-resolution Imagenette dataset [19] with atypical examples for the forget set. These results are primarily consistent with our findings on both CIFAR-10 as well as CIFAR-100 datasets, where models learned with our proposed unlearning algorithm are much more robust against relearning attacks.
All our prior evaluations focused on small resolution images, i.e., $32\times 32$ for both CIFAR-10 and CIFAR-100. In order to understand the generalizability of our approaches to higher resolution images, we evaluate on the higher-resolution ImageNette dataset [19], which is a subset of the ImageNet dataset [35]. The results are presented in Table˜1. Similarly to our findings on both CIFAR-10 (Section˜5) and CIFAR-100 (Appendix˜C), we see that the accuracy of retrain from scratch on the forget set after relearning stays close to 50%, while the accuracy of unlearning algorithms is significantly higher, except for our proposed weight dist reg algorithm among the tested methods.
Appendix E Tamper-Resistance of a Larger Model
@@
All our prior results focused on the smaller model, i.e., ResNet-18 [17]. In order to understand the impact of model size, we evaluate the susceptibility to relearning attacks on ResNet-34 (with almost double the number of parameters compared to ResNet-18, going from $\sim 11M$ to $\sim 21M$). The results are visualized in Fig.˜10. Note that we directly transfer the hyperparameter settings from our ResNet-18 experiments. Hence, we found that Weight Dist Reg, which was the most competitive method in terms of robustness, became susceptible to relearning. It is worth noting that the test set accuracy as well as the forget set accuracy for Weight Dist Reg in this case deviates significantly from all prior results, where it demonstrated consistently lower accuracies, highlighting a potential deficiency of the selected hyperparameters. Catastrophic Forgetting and Weight Dist Reg exhibit high robustness against relearning, which are instantiations of our framework.
Appendix F Evolution of Forget Set Accuracy During Relearning
Figure 11: Comparison between relearning time and accuracy on the held-out part of the forget set $\mathcal{D}_{F_{ho}}$ on CIFAR-10 and ResNet-18, where the forget set is comprised of atypical examples from the ‘airplane’ class. The figure indicates that many methods achieve near-perfect recovery of unlearned knowledge with only a small amount of model fine-tuning, without even assuming access to the unlearned examples. This figure is an extension of the results presented in Fig.˜2 (top).
Figure 12: Comparison between test set accuracy and accuracy on the held-out part of the forget set $\mathcal{D}_{F_{ho}}$ on CIFAR-10 and ResNet-18, where the forget set is comprised of atypical examples from all classes. This figure is an extension of the results presented in Fig.˜2 (bottom). We find that atypical examples, when selected from all the classes, are significantly harder in comparison to examples selected from a particular class, and hence, achieve a better separation between different methods in comparison to sub-class unlearning in Fig.˜2 (top).
In order to complement the results in Fig.˜2 (top) and Fig.˜3, we visualize the forget set accuracy evolution in Fig.˜11. We see that methods that are most stable are nearly unaffected by increasing training time, highlighting that more training time might not be sufficient to increase their susceptibility further.
Similarly, Fig.˜12 complements the results on class-agnostic unlearning in Fig.˜2 (bottom), where we visualize both the line plots as well as the results from two-phase training strategies. We visually observe best separation between different methods on the line plot in this case of class-agnostic unlearning.
Appendix G Relearning with Examples from the Same Distribution, which are Distinct from the Retain Set
@@
The results are presented in Fig.˜13. We observe that using the retain set $\mathcal{D}_{R}$ achieves higher accuracy on the forget set compared to other choices. However, these differences diminish as the number of relearning examples increases, since these examples directly represent the unlearned knowledge and may become the dominant factor in recovery. When the test set is used for relearning (instead of the retain set), we unsurprisingly observe perfect accuracy on the test set, as it serves as the training set in this scenario. These results indicate that relearning is more effective when using examples that were seen during training, rather than a held-out set of examples.
@@
Appendix H Membership-Inference Attacks (MIA) on Unlearned Models
Unlearning Method MIA Acc (%)
Retrain from Scratch 52.2%
L1 Sparse 54.2%
SCRUB 51.6%
Circuit Breakers 55.0%
Gradient Ascent 54.2%
Random Relabeling 94.8%
Catastrophic Forgetting 54.2%
SSD 53.8%
Weight Attenuation 53.6%
Weight Distortion 53.2%
Weight Dropout 54.8%
Weight Dist Reg 48.4%
Table 2: Membership-Inference Attack (MIA) accuracy, where 50% indicates chance-level performance. Note that most of the methods are close to chance-level performance, except random relabeling. This indicates that MIA attacks are not sufficient to understand susceptibility to relearning attacks.
Membership-Inference Attacks (MIA) attempt to evaluate if it is possible to identify instances that were part of the model’s training distribution from held-out instances [37]. Being able to distinguish instances leaks private information regarding an instance’s membership. We follow the setup from [26] for membership-inference attacks with balanced loss-threshold-based classifier. The results are presented in Table˜2. We observe that, as is the case in our initial findings with the forget set accuracy metric, this particular type of MIA can give a false sense that unlearning successfully erased all traces of unlearned data, even in cases where we know that the unlearned model is prone to relearning attacks. Note that we evaluated a particularly simple variant of MIA based on prior work [26]. Therefore, it is possible to use more complex variants of MIA, which might be better able to predict this vulnerability against relearning.
Appendix I Understanding the Efficiency of Unlearning Algorithms
Model Method Unlearning Epochs Test Acc (%) Forget Set Acc (%)
Unlearned Catastrophic Forgetting 1 93.79% 100.00%
Unlearned Catastrophic Forgetting 3 93.42% 100.00%
Unlearned Catastrophic Forgetting 10 92.52% 94.50%
Unlearned Catastrophic Forgetting 30 91.68% 77.75%
Unlearned Catastrophic Forgetting 50 90.25% 60.00%
Unlearned Catastrophic Forgetting 100 88.23% 41.75%
Unlearned Weight Distortion 1 84.49% 43.50%
Unlearned Weight Distortion 3 86.21% 49.00%
Unlearned Weight Distortion 10 86.27% 52.75%
Unlearned Weight Distortion 30 86.29% 53.25%
Unlearned Weight Distortion 50 86.31% 50.50%
Unlearned Weight Distortion 100 85.97% 45.75%
0 Relearn Catastrophic Forgetting 1 93.71% 99.99%
0 Relearn Catastrophic Forgetting 3 93.65% 99.99%
0 Relearn Catastrophic Forgetting 10 93.37% 99.28%
0 Relearn Catastrophic Forgetting 30 93.22% 93.63%
0 Relearn Catastrophic Forgetting 50 93.08% 82.16%
0 Relearn Catastrophic Forgetting 100 92.80% 66.14%
0 Relearn Weight Distortion 1 90.14% 61.07%
0 Relearn Weight Distortion 3 90.14% 61.66%
0 Relearn Weight Distortion 10 89.86% 61.38%
0 Relearn Weight Distortion 30 89.62% 60.69%
0 Relearn Weight Distortion 50 89.44% 59.64%
0 Relearn Weight Distortion 100 89.64% 58.59%
Table 3: Test and forget-set accuracy of the unlearned model after relearning attack on CIFAR-10 with atypical forget set examples. It is evident from the results that even a small number of unlearning epochs are sufficient to achieve robustness against relearning attack with weight distortion. In contrast, other methods such as catastrophic forgetting only work at the largest epoch setting. Note that all previous results were reported with 100 unlearning epochs, masking any impact of the efficiency of the unlearning algorithm.
In order to control for efficiency effects of the unlearning algorithm, all previous results were reported at a fixed unlearning budget of 100 epochs, which is comparable to the budget allocated for pretraining. This raises concerns regarding the utility of approximate unlearning algorithms, as one should use the gold-standard unlearning algorithm, i.e., retrain from scratch in cases where the computational budget is directly comparable to the pretraining budget.
In order to understand the significance of this selection, we vary the number of unlearning epochs and report test accuracy and forget set accuracy after relearning, while using the same set of hyperparameters. The results are presented in Table˜3. It is clear from the table that while both catastrophic forgetting and weight distortion are successful in resisting relearning at the maximum budget of 100 epochs, catastrophic forgetting is still susceptible to relearning with a smaller number of unlearning epochs. On the other hand, weight distortion is more resistant to relearning even after a single epoch of unlearning.
Appendix J Computational Resources